I dont know how to do the last part? could someone help meIn a survey of 160 randomly sampled households, 72 report owning a dog. Find a 99% confidence interval for the proportion of
Answer #1

of Given In a Lurney households 72 report 160 owing randomly sampled a dogan 3 na 160 d F = 72 x = 72 - 0.45 160 so ? = 1-0.4

answered by: ANURANJAN SARSAM
Answer #2

of Given In a Lurney households 72 report 160 owing randomly sampled a dogan 3 na 160 d F = 72 x = 72 - 0.45 160 so ? = 1-0.4